OpenStudy (anonymous):

simplify the expression and write in standard form square root od -144+3

OpenStudy (sonic):

\[\sqrt {-141}=\sqrt{-1}*\sqrt{141}=i \sqrt{141}\]

OpenStudy (sonic):

47*3=141. \[\sqrt{141}\] cannot be simplified.